# Supernova Electron-Neutrino Interactions with Xenon in the nEXO Detector

### v1

@@ IV.3 Event formation
After reconstructing light and charge, these events were correlated with each other. The approach taken was to form a list of all possible charge clusters that reconstruct to valid positions within the TPC when assuming correlation with the prompt light signal, and to do the same for the delayed light signal (due to neutron capture) if present. A charge cluster which reconstructed a valid interaction position for only one of the light signals was assumed to be correctly associated and eliminated as a candidate charge event for the other light signal. If degeneracies were still present, the largest charge signal was assumed to be correlated with the largest light signal, which has an impact on the reconstructed energy resolution.
Using the location of the individual depositions within a charge cluster, and a map of the geometric dependence of nEXO’s photon collection efficiencies, a weighted-average light collection efficiency was determined for each event proportional to the charge at each x- and y- channel intersection point. This efficiency is used to estimate the true number of photons produced by the scintillation signal, and thus the calibrated energy of the light signal.
This example algorithm was found to reconstruct electron-neutrino interactions with 88.4% efficiency. The remaining events were either missing a charge signal (no depositions within the drift region or above the charge-channel threshold) or a light signal (no depositions above scintillation-light threshold). While our reconstruction algorithm is not able to obtain an energy estimate for this $\sim$11.6of events, it may still be possible to use them to further understand supernova electron-neutrino neutrino interactions.
@@ V Reconstructing the supernova neutrino spectrum
Neutrinos of different energies and flavors decouple from thermal equilibrium at different depths in a CCSN. Thus, reconstructing the neutrino energy spectrum of different flavors allow for imaging of the interior dynamics of the collapse. Many factors can impact a detector’s ability to reconstruct the incident neutrino spectrum, such as statistics, energy resolution, and detection threshold. There are a number of existing studies [37, 14, 73, 74, 75, 76] evaluating the ability of different detectors to reconstruct an incident neutrino spectrum given an observed detector energy spectrum, but none of these focus on xenon-detectors using the charged-current channel.
@@ V.1 Pinched-thermal spectrum
A complete description of the expected supernova neutrino signal is complex, as it is affected by many factors, such as the neutrino mass splitting, self-induced flavor conversion, progenitor mass, and the MSW (Mikheyev–Smirnov–Wolfenstein) effect [77]. While a complete description of the expected neutrino spectrum from a CCSN is beyond the scope of this study, more detail on the impact of these factors on the neutrino spectrum can be found in Refs. [6, 78]. Here we adopt a simplified pinched-thermal neutrino spectrum [79, 75],
@@
where $d$ is the distance to the collapsing star, $\varepsilon$ is the total binding energy released in the form of neutrinos by the CCSN, $\alpha$ is a pinching parameter representing the deviation from a Fermi-Dirac spectrum, and $\langle E_{\nu}\rangle$ is the mean energy of the neutrinos. While there are many sets of parameters that may describe the spectrum of electron neutrinos incident on an earth-based detector integrated over the $\sim$10-second CCSN burst, we use the same example parameters as in Ref. [14] to allow for a comparison of the results within that reference. The example CCSN parameters chosen in that references are $(\alpha,\varepsilon,\langle E_{\nu}\rangle)=(2.5,5\times 10^{52}\mbox{ ergs},9.5\mbox{ MeV})$.
There are a variety of other models formulated to describe the expected neutrino spectrum (such as Refs. [50, 51, 80]), but here we focus on only the pinched-thermal form as it is easy to parameterize. In principle, this work could be extended to other models using the transfer matrix we develop (Sec. V.2). We also assume the distance to the collapsing star is well-known, and fix this distance parameter in describing the spectrum. While not of relevance for this study, the electron-antineutrino component and NC component ($\nu_{\mu}+\nu_{\tau}+\bar{\nu_{\mu}}+\bar{\nu_{\tau}}$) can also be described by a pinched-thermal form, with a different set of parameters. Fig. 5 depicts these three components with parameter assumptions from Ref. [14].
Figure 5: Pinched-thermal spectrum with $(\alpha,\varepsilon,\langle E_{\nu}\rangle)=(2.5,5\times 10^{52}\mbox{ ergs},9.5\mbox{ MeV})$ for $\nu_{e}$, $(2.5,5\times 10^{52}\mbox{ ergs},12\mbox{ MeV})$ for $\bar{\nu_{e}}$, and $(2.5,5\times 10^{52}\mbox{ ergs},15.6\mbox{ MeV})$ for each flavor in the $\nu_{x}$ component.
@@ V.2 Transfer matrix
Using the methods described in the above text, one million $\nu_{e}-{}^{134}$eCC, $\nu_{e}-{}^{136}$eCC, and $\nu_{e}-e^{-}$ were each generated with MARLEY spanning a uniform energy distribution from 0 to 100 MeV, simulated using the nEXO simulation framework, and analyzed with the previously-discussed reconstruction algorithm. A transfer matrix was formed for each interaction type mapping incident neutrino energy into reconstructed detector energy. In the analysis, the individual transfer matrices from these three components were used, but to illustrate the total detector response these are combined with cross section and isotopic abundance weighting in Fig. 6. The individual transfer matrices for each component can be found in the appendix.
@@
While cuts on the data can improve the precision with which nEXO reconstructs the true energy deposited by a particle, leading to improved energy resolution, a minimal set of cuts were employed here to maximize statistics. A cut was developed to remove events with a reconstructed charge deposition center within 5 cm of the nEXO cathode. This is greater than the standoff cut employed in the $0\nu\beta\beta$ analysis in part due to the large spatial size of these events leading to more energy deposited below the cathode where light collection efficiency is worse. Adopting this cut reduces the overall reconstruction efficiency from 88.4% to 86.0%. The transfer matrix utilized preserves the energy-dependence and overall normalization of the reconstruction efficiency.
@@ V.3 Fitting procedure
The neutrino interaction events generated by MARLEY were run through the nEXO simulation framework and the described reconstruction algorithm. Using the reconstructed spectrum from the nominal set of pinched-thermal spectrum parameters, 1.5 million data sets (referred to as test data sets) were generated with Poisson statistics for supernovae located at distances of 0.2 and 0.5 kpc from Earth.
Using RooFit [81], a likelihood function was developed to compare the simulated energy spectrum in the detector using a trial set of pinched-thermal parameters to the simulated energy spectrum of the test data set. First a trial pinched-thermal neutrino spectrum was generated and multiplied by the predicted cross sections from MARLEY for $\nu_{e}-{}^{134}$eCC, $\nu_{e}-{}^{136}$eCC, and $\nu_{e}-e^{-}$ interactions. Next, the xenon mass and isotopic abundance of nEXO were used to generate the expected number of neutrino interactions of each type. Then, the transfer matrices were used to incorporate detector response and generate extended probability distribution functions (PDFs) for each type of interaction, which were combined into a single extended PDF. The combined PDF had a bin-size of 0.2 MeV, and the test data sets fit to were unbinned. The likelihood function was then minimized using the iminuit python package [82].
@@ V.4 Fit results
Following each fit to a test data set, the difference between the minimum negative log-likelihood ($LL_{\mbox{min}}$) and the negative log-likelihood calculated using the nominal set of parameters ($LL_{\mbox{true}}$) was determined, referred to as $-\Delta LL=-(LL_{\mbox{min}}-LL_{\mbox{true}})$. Due to the observed non-Gaussian nature of the distribution of $-\Delta LL$ values, one- and two-sigma confidence intervals were generated numerically by selecting $-\Delta LL$ values encapsulating 68.2% and 95.5% of the data sets, respectively. These bands are plotted for data sets generated at 0.2 and 0.5 kpc in Fig. 7, using the corner [83] package. Black lines indicate the nominal parameters of the incident supernova neutrino spectrum. The 1D histograms show the complete distribution of the posterior probability distributions for each parameter.
@@
Figure 7: The two-dimensional Gaussian one- and two-sigma contours are drawn for the posterior probability distributions for CCSN originating 0.2 and 0.5 kpc from Earth, utilizing 1.5 million data sets generated with Poisson statistics. The nominal set of parameters $(\alpha,\varepsilon,\langle E_{\nu}\rangle)$ of (2.5, 5×$\times 10^{52}$ ergs, 9.5 MeV) are indicated by the black lines.
@@
Figure 8: One- and two-sigma confidence intervals showing nEXO’s ability to reconstruct incident electron neutrino parameters as a function of distance. The dashed black line indicates the true parameter values, and the vertical gray lines indicate the location of nearby CCSN candidates, from Ref. [52].
@@
To study nEXO’s ability to reconstruct incident electron neutrino parameters as a function of CCSN distance, 60k data sets are generated with a step size of 0.1 kpc. One- and two-sigma confidence intervals are selected based on the $-\Delta LL$ values enclosing 68.2% and 95.5% of the data, and are plotted in Fig. 8 along with the true values. The gray vertical lines show the locations of nearby RSGs, from Ref. [52]. At greater distances, many of the data sets do not have a sufficient number of events to capture the high-energy tail of the pinched-thermal distribution, resulting in a large increase in the spread of the $\alpha$ parameter. While limited statistics impact nEXO’s ability to reconstruct the incident neutrino spectrum parameters, data from nEXO could be combined with that from other detectors to further constrain these parameters when statistics are poor.
@@
Figure 9: Distribution of posterior probability distributions for parameters are shown in the one-dimensional histograms, along with median values of the distribution, for a 300-tonne liquid xenon detector (90% 136Xe, 10% 134Xe) reconstructing events from a supernovae located 4 kpc away. The two-dimensional plot shows 90% confidence intervals for this detector, as well as for a 40-kilotonne liquid argon detector, from Ref. [14].
@@
These results are extended to a 300-tonne liquid xenon detector based of the same isotopic abundance using the same transfer matrix. While this is an approximation, as the ratio of xenon-skin to TPC-volume will likely be different for a 300-tonne detector, Fig. 9 illustrates how a detector of this size would compare to the 40-kilotonne liquid argon detector considered in Ref. [14]. While there are a number of challenges that need to be addressed to develop a xenon-detector of this size [18], with its increased mass compared to that of nEXO, a 300-tonne detector can reconstruct the incident neutrino spectrum to much further distances, and reconstruct nearby supernova electron-neutrino parameters with improved accuracy.
@@ V.5 Future modeling improvements
The findings of this study can potentially be impacted by a number of factors whose contributions are difficult to quantify. The cross section and its energy-dependence can be impacted by the value of the nuclear matrix elements and contributions from forbidden transitions not accounted for in the MARLEY model. While the calculation using MARLEY shows a general agreement with the calculations in Ref. [22], uncertainties could be better quantified if experimental data existed for electron neutrinos in a similar energy range. The electron-neutrino spectrum shape and normalization impact the expected number of events that would be observed in nEXO. While only a handful of supernova neutrino events have been previously observed [84, 85, 86], future supernova neutrino interactions would test these models. Although uncertainties on the spins, parities, and branching ratios of excited states of 136Cs are not expected to have a large impact on the visible spectrum from supernova neutrino interactions, experimental measurements of these quantities would lead to higher fidelity simulations. Finally, there may additionally be energy deposited from electron-antineutrino interactions and NC interactions which can affect the ability to reconstruct the true electron-neutrino parameters, although these contributions are also expected to be small [49, 22].
@@ VI Conclusion
There is a wealth of physics that can be unlocked by studying the neutrinos from the next nearby core-collapse supernova, provided there are detectors capable of observing these interactions whose response is well-known. We have evaluated the sensitivity of the nEXO detector to electron-neutrino interactions from supernovae using the MARLEY event generator, simulating CC interactions on the 134Xe and 136Xe isotopes. Events generated by MARLEY have been simulated using the nEXO simulation framework, and reconstructed using a custom algorithm designed to accommodate the unique nature of CC events compared to the signal and background events of primary interest to nEXO. Based on our work, we estimate that nEXO will be able to observe electron-neutrino interactions with xenon from supernovae as far as 5 to 8 kpc from earth, depending on the supernova neutrino flux model used and parameters of the collapsing star.
A transfer matrix was developed to map incident electron neutrino energy into observed energy in nEXO, which is used to reconstruct the incident supernova neutrino spectrum using a pinched-thermal parameterization. It is observed that uncertainties in the parameter reconstruction grow rapidly for distances above $\sim$0.5kpc, where insufficient statistics in the tail of the spectrum limit the ability to reconstruct the pinching parameter.
These results were extended to a 300-tonne xenon detector of the same isotopic abundance, which is able to detect supernova electron neutrinos out to further distances, covering most of the candidates listed within a recent survey of RSGs. While a 300-tonne detector would not be able to reconstruct supernova neutrino parameters with as much precision as DUNE, it would still provide useful information about the electron-neutrino component of nearby CCSN and contribute to the global dataset of all supernova neutrino detectors.

### v2

@@ IV.3 Event formation
After reconstructing light and charge, these events were correlated with each other. The approach taken was to form a list of all possible charge clusters that reconstruct to valid positions within the TPC when assuming correlation with the prompt light signal and to do the same for the delayed light signal (due to neutron capture) if present. A charge cluster that reconstructed a valid interaction position for only one of the light signals was assumed to be correctly associated and eliminated as a candidate charge event for the other light signal. If degeneracies were still present, the largest charge signal was assumed to be correlated with the largest light signal, which has an impact on the reconstructed energy resolution.
Using the location of the individual depositions within a charge cluster and a map of the geometric dependence of nEXO’s photon collection efficiencies, a weighted-average light collection efficiency was determined for each event proportional to the charge at each x- and y-channel intersection point. This efficiency is used to estimate the true number of photons produced by the scintillation signal and thus the calibrated energy of the light signal.
This example algorithm was found to reconstruct electron-neutrino interactions generated with a pinched-thermal spectrum described above with 88.4% efficiency. The remaining events were either missing a charge signal (no depositions within the drift region or above the charge-channel threshold) or a light signal (no depositions above scintillation light threshold). While our reconstruction algorithm is not able to obtain an energy estimate for this $\sim$11.6of events, it may still be possible to use them to further understand supernova electron-neutrino neutrino interactions.
Using the reconstructed events, we can tag neutron captures by looking for the presence of multiple light signals separated in time by at least 300 ns. Neutron captures were observed for $\sim$33.6of the true neutron-emitting events. The identification of neutron-emitting events could potentially be improved by developing an algorithm based on event topology or using machine learning.
@@ V Reconstructing the supernova neutrino spectrum
Neutrinos of different energies and flavors decouple from thermal equilibrium at different depths in a CCSN. Thus, reconstructing the neutrino energy spectrum of different flavors allow for imaging of the interior dynamics of the collapse. Many factors can impact a detector’s ability to reconstruct the incident neutrino spectrum, such as statistics, energy resolution, and detection threshold. There are a number of existing studies [33, 14, 75, 76, 77, 78] evaluating the ability of different detectors to reconstruct an incident neutrino spectrum given an observed detector energy spectrum, but none of these focus on xenon-based detectors using the charged-current channel.
@@ V.1 Pinched-thermal spectrum
A complete description of the expected supernova neutrino signal is complex, as it is affected by many factors, such as the neutrino mass splitting, self-induced flavor conversion, progenitor mass, and the Mikheyev–Smirnov–Wolfenstein effect [79]. While a complete description of the expected neutrino spectrum from a CCSN is beyond the scope of this study, more detail on the impact of these factors on the neutrino spectrum can be found in Refs. [6, 80]. Here we adopt a simplified pinched-thermal neutrino spectrum [81, 77],
@@
where $d$ is the distance to the collapsing star, $\varepsilon$ is the total binding energy released in the form of neutrinos by the CCSN, $\alpha$ is a positive-definite pinching parameter representing the deviation from a Fermi-Dirac spectrum, and $\langle E_{\nu}\rangle$ is the mean energy of the neutrinos. While there are many sets of parameters that may describe the spectrum of electron neutrinos incident on an Earth-based detector integrated over the $\sim$10-sec CCSN burst, we use the same example parameters as in Ref. [14] to allow for a comparison of the results within that reference. The example CCSN parameters chosen in that reference are $(\alpha,\varepsilon,\langle E_{\nu}\rangle)=(2.5,5\times 10^{52}\mbox{ erg},9.5\mbox{ MeV})$.
There are a variety of other models formulated to describe the expected neutrino spectrum (such as Refs. [50, 51, 82]), but here we focus on only the pinched-thermal form as it is easy to parameterize. In principle, this work could be extended to other models using the transfer matrix we develop (Sec. V.2). We also assume the distance to the collapsing star is well known, and fix this distance parameter in describing the spectrum. While not of relevance for this study, the electron-antineutrino component and NC component ($\nu_{\mu}+\nu_{\tau}+\bar{\nu_{\mu}}+\bar{\nu_{\tau}}$) can also be described by a pinched-thermal form, with a different set of parameters. Figure 5 depicts these three components with parameter assumptions from Ref. [14].
Figure 5: Pinched-thermal spectrum with $(\alpha,\varepsilon,\langle E_{\nu}\rangle)=(2.5,5\times 10^{52}\mbox{ erg},9.5\mbox{ MeV})$ for $\nu_{e}$, $(2.5,5\times 10^{52}\mbox{ erg},12\mbox{ MeV})$ for $\bar{\nu_{e}}$, and $(2.5,5\times 10^{52}\mbox{ erg},15.6\mbox{ MeV})$ for each flavor in the $\nu_{x}$ component.
@@ V.2 Transfer matrix
Using the methods described in the above text, $1\times 10^{6}$ $\nu_{e}-{}^{134}$eCC, $\nu_{e}-{}^{136}$eCC, and $\nu_{e}-e^{-}$ were each generated with MARLEY spanning a uniform energy distribution from 0 to 100 MeV, simulated using the nEXO simulation framework, and analyzed with the previously discussed reconstruction algorithm. A transfer matrix was formed for each interaction type mapping incident neutrino energy into reconstructed detector energy. In the analysis, the individual transfer matrices from these three components were used, but to illustrate the total detector response these are combined with cross section and isotopic abundance weighting in Fig. 6. The individual transfer matrices for each component can be found in Appendix C.
@@
While cuts on the data can improve the precision with which nEXO reconstructs the true energy deposited by a particle, leading to improved energy resolution, a minimal set of cuts were employed here to maximize statistics. A cut was developed to remove events with a reconstructed charge deposition center within 5 cm of the nEXO cathode. This is greater than the standoff cut employed in the $0\nu\beta\beta$ analysis in part due to the large spatial size of these events leading to more energy deposited below the cathode where light collection efficiency is worse. Adopting this cut reduces the overall reconstruction efficiency from 88.4% to 86.0%. The transfer matrix utilized preserves the energy dependence and overall normalization of the reconstruction efficiency.
@@ V.3 Fitting procedure
The neutrino interaction events generated by MARLEY were run through the nEXO simulation framework and the described reconstruction algorithm. Using the reconstructed spectrum from the nominal set of pinched-thermal spectrum parameters, $1.5\times 10^{6}$ datasets (referred to as test datasets) were generated with Poisson statistics for supernovae located at distances of 0.2 and 0.5 kpc from Earth.
Using RooFit [83], a likelihood function was developed to compare the simulated energy spectrum in the detector using a trial set of pinched-thermal parameters to the simulated energy spectrum of the test dataset. First, a trial pinched-thermal neutrino spectrum was generated and multiplied by the predicted cross sections from MARLEY for $\nu_{e}-{}^{134}$eCC, $\nu_{e}-{}^{136}$eCC, and $\nu_{e}-e^{-}$ interactions. Next, the xenon mass and isotopic abundance of nEXO were used to generate the expected number of neutrino interactions of each type. Then, the transfer matrices were used to incorporate detector response and generate extended probability distribution functions (PDFs) for each type of interaction, which were combined into a single extended PDF. The combined PDF had a bin size of 0.2 MeV, and the test datasets fit to were unbinned. The likelihood function was then minimized using the iminuit python package [84].
@@ V.4 Fit results
Following each fit to a test dataset, the difference between the minimum negative log-likelihood ($LL_{\mbox{min}}$) and the negative log-likelihood calculated using the nominal set of parameters ($LL_{\mbox{true}}$) was determined, referred to as $-\Delta LL=-(LL_{\mbox{min}}-LL_{\mbox{true}})$. Because of the observed non-Gaussian nature of the distribution of $-\Delta LL$ values, one- and two-$\sigma$ confidence intervals (C.I.s) were generated by profiling over the nuisance parameters and selecting $-\Delta LL$ values that encapsulate 68.2% and 95.5% of the datasets, as shown in Fig.7. These bands are plotted for datasets generated at 0.2 and 0.5,kpc using the corner[85] package. Black lines indicate the nominal parameters of the incident supernova neutrino spectrum. The 1D histograms show the posterior probability distributions for each parameter.
@@
Figure 7: One- and two-sigma contours are drawn for the posterior probability distributions for CCSNe originating 0.2 and 0.5 kpc from Earth, utilizing $1\times 10^{6}$ datasets generated with Poisson statistics. The nominal set of parameters $(\alpha,\varepsilon,\langle E_{\nu}\rangle)$ of (2.5, 5×$\times 10^{52}$ erg, 9.5 MeV) are indicated by the black lines.
@@
Figure 8: One- and two-$\sigma$ confidence intervals showing nEXO’s ability to reconstruct incident electron-neutrino parameters as a function of distance. The dashed black line indicates the true parameter values, and the vertical gray lines indicate the location of nearby CCSNe candidates, from Ref. [53].
@@
To study nEXO’s ability to reconstruct incident electron neutrino parameters as a function of CCSN distance, 60,000 datasets are generated with a step size of 0.1 kpc. One- and two-$\sigma$ confidence intervals are formed by profiling over the nuisance parameters and selecting $-\Delta LL$ values enclosing 68.2% and 95.5% of the data, which are plotted in Fig. 8 along with the true values of these parameters. The gray vertical lines show the locations of nearby RSGs, from Ref. [53]. At greater distances, many of the datasets do not have a sufficient number of events to capture the high-energy tail of the pinched-thermal distribution, resulting in a large increase in the spread of the $\alpha$ parameter. While limited statistics impact nEXO’s ability to reconstruct the incident neutrino spectrum parameters, data from nEXO could be combined with that from other detectors to further constrain these parameters when statistics are poor.
@@
Figure 9: Distribution of posterior probability distributions for parameters are shown in the one-dimensional histograms, along with median values of the distribution, for a 300-metric-ton liquid xenon detector (90% 136Xe, 10% 134Xe) reconstructing events from a supernova located 4 kpc away. The two-dimensional plot shows 90% confidence intervals for this detector, as well as for a 40-metric-kt liquid argon detector, from Ref. [14].
@@
These results are extended to a 300-metric-ton liquid xenon detector based on the same isotopic abundance using the same transfer matrix. While this is an approximation, as the ratio of xenon skin to TPC volume will likely be different for a 300-metric-ton detector, Fig. 9 illustrates how a detector of this size would compare to the 40-metric-kt liquid argon detector considered in Ref. [14]. While there are a number of challenges that need to be addressed to develop a xenon detector of this size [17], with its increased mass compared to that of nEXO, a 300-metric-ton detector can reconstruct the incident neutrino spectrum to much further distances, and reconstruct nearby supernova electron-neutrino parameters with improved accuracy.
@@ V.5 Future modeling improvements
The findings of this study can potentially be impacted by a number of factors whose contributions are difficult to quantify. The cross section and its energy dependence can be impacted by the value of the nuclear matrix elements and contributions from forbidden transitions not accounted for in the MARLEY model. While the calculation using MARLEY shows a general agreement with the calculations in Ref. [21], uncertainties could be better quantified if experimental data existed for electron neutrinos in a similar energy range. The electron-neutrino spectrum shape and normalization impact the expected number of events that would be observed in nEXO. While only a handful of supernova neutrino events have been previously observed [86, 87, 88], future supernova neutrino interactions would test these models. Although uncertainties on the spins, parities, and branching ratios of excited states of 136Cs are not expected to have a large impact on the visible spectrum from supernova neutrino interactions, experimental measurements of these quantities would lead to higher fidelity simulations. Finally, there may be additional energy deposited from electron-antineutrino interactions and NC interactions that can affect the ability to reconstruct the true electron-neutrino parameters, although these contributions are also expected to be small [46, 21].
@@ VI Conclusion
There is a wealth of physics that can be unlocked by studying the neutrinos from the next nearby core-collapse supernova, provided there are detectors capable of observing these interactions whose response is well known. We have evaluated the sensitivity of the nEXO detector to electron-neutrino interactions from supernovae using the MARLEY event generator, simulating CC interactions on the 134Xe and 136Xe isotopes. Events generated by MARLEY have been simulated using the nEXO simulation framework, and reconstructed using a custom algorithm designed to accommodate the unique nature of CC events compared to the signal and background events of primary interest to nEXO. Based on our work, we estimate that nEXO will be able to observe electron-neutrino interactions with xenon from supernovae as far as 5-8 kpc from Earth, depending on the supernova neutrino flux model used and parameters of the collapsing star.
A transfer matrix was developed to map incident electron-neutrino energy into observed energy in nEXO, which is used to reconstruct the incident supernova neutrino spectrum using a pinched-thermal parameterization. We find that uncertainties in the parameter reconstruction grow rapidly for distances above $\sim$0.5kpc, where insufficient statistics in the tail of the spectrum limit the ability to reconstruct the pinching parameter.
These results were extended to a 300-metric-ton xenon detector of the same isotopic abundance, which is able to detect supernova electron neutrinos out to further distances, covering most of the candidates listed within a recent survey of RSGs. While a 300-metric-ton detector would not be able to reconstruct supernova neutrino parameters with as much precision as DUNE, it would still provide useful information about the electron-neutrino component of nearby CCSNe and contribute to the global dataset of all supernova neutrino detectors.